OpenStudy (anonymous):

solve the equation y=2x for x solve the equation 3y +z = 13 for z how do you do these problems? please explain along with the answer

OpenStudy (triciaal):

y = 2x what do you have to do to 2x to get 1x?

OpenStudy (anonymous):

you 2x by 2

OpenStudy (anonymous):

divide*

OpenStudy (triciaal):

correct what you do to one side of an equation you do to the other to keep the balance so since you divide by 2 on the right you do the same on the left

OpenStudy (anonymous):

oh, so it would by y over 2? thanks for the help!

OpenStudy (triciaal):

agree x = y/2?

OpenStudy (triciaal):

for the 2nd one same principle isolate the variable to one side of the equation then do what is necessary to make the coefficient = 1 and remember to keep the balance.

OpenStudy (triciaal):

3y +z = 13 to isolate z to one side what do you need to do?

OpenStudy (anonymous):

i subtract the 3y would that mean that x=13-3y?

OpenStudy (triciaal):

z

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

alright thanks!

OpenStudy (triciaal):

welcome